OpenStudy (anonymous):

A school system is reducing the amount of dumpster loads of trash removed each week. In week 5, there were 45 dumpster loads of waste removed. In week 10, there were 30 dumpster loads removed. Assume that the reduction in the amount of waste each week is linear. Write an equation in function form to show the amount of trash removed each week. f(x) = - 3x + 45 f(x) = 3x + 45 f(x) = - 3x + 60 f(x) = 3x + 60

OpenStudy (anonymous):

Somebody?

OpenStudy (anonymous):

I will give medal

OpenStudy (anonymous):

y = -3x + 60 { {5, 45}, {6, 42}, {7, 39}, {8, 36}, {9, 33}, {10, 30} }

OpenStudy (anonymous):

so what does that mean @robtobey

OpenStudy (anonymous):

heeelp @pooja195

OpenStudy (anonymous):

f(x) is the line with the RHS equal to -3x+60 . The second line is a list of the points where x=5 through x=10 in steps of one.

OpenStudy (anonymous):

y = f(x).

OpenStudy (anonymous):

@thefurball @SolomonZelman i don't understand this

OpenStudy (anonymous):

@Aureyliant @pooja195 @misssunshinexxoxo

OpenStudy (misssunshinexxoxo):

y-30=-3(x-10) y-30=-3x+30 y=-3x+60 Answer C

OpenStudy (misssunshinexxoxo):

Cheers dragon! :)

OpenStudy (anonymous):

thank you!!! @misssunshinexxoxo